TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Dan Dubrick
date: 2003-04-15 23:40:00
subject: 4\04 Pt 2 ISS On Orbit Status 04-04-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

05 Apr 2003

ISS On-Orbit Status 4 Apr 2003

Part  2 of 2

Today's CEO (crew earth observations) targets, currently restricted
by flight rule constraints limiting the use of the science window in
the Lab, were East Asian smoke (Dynamic event:  Fires in eastern
Russia are giving rise to a smoke pall between 40N and 50N north of
Japan.  Shooting obliques left and right of track), Eastern
Mediterranean Dust (dust blowing off North Africa yesterday should
continue today), Lake Poopo, Bolivia (good nadir pass to document
lake shorelines), and Smoke in the southern US (Dynamic event:
Looking left and right for fires and smoke in Texas, Louisiana,
Arkansas and Missouri, shooting obliques). CEO images can be viewed
at the websites http://eol.jsc.nasa.gov and
http://earthobservatory.nasa.gov

U.S. and Russian Segment Status (as of 12:45pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control
(TCS): Elektron O2 generator is powered On.  Vozdukh CO2 scrubber is
On (manual mode 5).  U.S. CDRA CO2 scrubber is Off.  TCCS (trace
contaminant control subsystem) is operating.  MCA (major constituents
analyzer) is operating.  BMP Harmful Impurities unit: Absorbent bed
#1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1
is On; SKV-2 is Off.

SM Working Compartment:  Pressure (mmHg) -- 738; temperature (deg
C) -- 27.4; ppO2 (mmHg) -- 183.1 (suspect); ppCO2 (mmHg) -- 2.3 (data
suspect). SM Transfer Compartment:  Pressure (mmHg) -- 741;
temperature (deg C) -- 19.8.

FGB Cabin:  Pressure (mmHg) -- 755; temperature (deg C) -- 19.8.
Node:  Pressure (mmHg) -- 736.89; temperature (deg C) -- 23.8
(shell); ppO2 (mmHg) -- 165.1; ppCO2 (mmHg) -- 4.0.

U.S. Lab:  Pressure (mmHg) -- 739.16; temperature (deg C) -- 22.3;
ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) -- 739.16; temperature
(deg C) -- 20.8; shell heater temp (deg C) -- 21.0, ppO2 (mmHg) --
n/a; ppCO2 (mmHg) -- n/a.
PMA-1:  Shell heater temp (deg C) -- 23.7
PMA-2:  Shell heater temp (deg C) -- 19.7

(n/a = data not available)

Propulsion System (PS):

Total propellant load available [SM(774) + FGB(2746) + Progress(474]
--- 3994 kg (8805 lb) as of 4/3/03.  (Capability: SM -- 860 kg;
FGB -- 6120 kg). 

Electrical Power Systems (EPS):

Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B
and 4B both in dual-angle ("blind") mode.
SM batteries:  All batteries (8) are in "Partial Charge" mode.
FGB batteries:  Battery #5 is disconnected; battery #3 is in "Cycle"
mode; all other batteries (5) are in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 and PCU-2 both in Standby mode.

Command & Data Handling Systems:

C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
GNC-2 MDM is prime; GNC-1 is Backup.
INT-1 is operating; INT-2 is Off.
EXT-1 is On (primary), EXT-2 is Off.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-2 MDM is On (primary); PL-1 MDM is Off
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
Being upgraded.
SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
Being upgraded.

Attitude Control Systems:

3 CMGs on-line (CMG-1 failed).
State vector source -- U.S. SIGI-1 (GPS)
Attitude source -- U.S. SIGI-1 (GPS)
Angular rate source -- RGA-2

Flight Attitude:

XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: -179.0
deg, pitch: -8.8 deg., roll: 0 deg]), with CMG TA (thruster assist)
Momentum Management.
Solar Beta Angle:  16.6 deg (magnitude decreasing).

Communications & Tracking Systems:

FGB MDM-1 is powered Off; FGB MDM-2 is operational.
All other Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally (VTR1 is operable again).
HCOR (high-rate communications outage recorder) is operating
nominally.

Robotics:

SSRMS/Canadarm2 based at Lab PDGF with Keep Alive (KA) power on both
strings.
MBS: KA power on both strings.
MT: latched at WS4, with KA power.
POA: KA power on both strings.
RWS (robotics workstations): Lab RWS is Off; Cupola RWS is Off.

 - END OF FILE -
==========

@Message posted automagically by IMTHINGS POST 1.30
--- 
* Origin: SpaceBase(tm) Pt 1 -14.4- Van BC Canada 604-473-9358 (1:153/719.1)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.